# Test-Time Scaling with Reflective Generative Model

### v1

@@ Abstract
We introduce our first reflective generative model MetaStone-S1, which obtains OpenAI o3’s performance via the self-supervised process reward model (SPRM). Through sharing the backbone network and using task-specific heads for next token prediction and process scoring respectively, SPRM successfully integrates the policy model and process reward model(PRM) into a unified interface without extra process annotation, reducing over 99% PRM parameters for efficient reasoning. Equipped with SPRM, MetaStone-S1 is naturally suitable for test time scaling (TTS), and we provide three reasoning effort modes (low, medium, and high), based on the controllable thinking length. Moreover, we empirically establish a scaling law that reveals the relationship between total thinking computation and TTS performance. Experiments demonstrate that our MetaStone-S1 achieves comparable performance to OpenAI-o3-mini’s series with only 32B parameter size. To support the research community, we have open-sourced MetaStone-S1 at https://github.com/MetaStone-AI/MetaStone-S1.
@@ 1 Introduction
TTS approaches can be categorized into two types: internal TTS and external TTS. Internal TTS strategies use CoT for longer thinking processes (Guo et al., 2025; OpenAI, 2024), which benefits from Long-CoT Supervised Fine-Tuning and reinforcement learning. Recent internal TTS methods (Guo et al., 2025) mainly suffer from the false positive reasoning process, as the outcome reward will misclassify the correct answer with incorrect reasoning during the training stage. External TTS is proposed for selecting correct reasoning process, which is proved to be more effective in performance boosting compared with outcome reward (Lightman et al., 2023). Prominent external TTS algorithms include Best-of-N sampling, Beam Search, and Diverse Verifier Tree Search, using the Process Reward Model(PRM) to select the sampling and search-based candidates. For example, Best-of-N firstly generates multiple outputs, and then sequentially uses PRM-based to score and select the best solution. Liu et al. (2025) point out that the best external TTS solution varies from policy model with different parameters. When the parameter of the policy model is less than 32B, the search methods (Beam Search and Diverse Verifier Tree Search) achieve better results than the sampling one(Best-of-N). However, when the parameter size is equal to or greater than 32B, the sampling method can achieve better performance. Based on the above analysis, internal and external TTS are two individual methods and can benefit each other, e.g. using the Best-of-N to boost the Long-CoT model with PRM.
This paper focuses on external TTS and proposes a Reflective Generative Form that enables LLMs to integrate reasoning and evaluation within a single model. We first design a Self-supervised Process Reward Model (SPRM), which shares its backbone network with the policy model. By employing task-specific heads for next-token prediction and process scoring, SPRM effectively unifies the policy model and process reward model(PRM) into a unified interface. Compared to previous PRMs, this unified design provides a more efficient scoring process with little parameter and computational overhead. To reduce the need for expensive process annotations, we further introduce a Self-supervised Process Reward Loss, which optimizes the SPRM using only final answers. Based on SPRM, we propose MetaStone-S1, containing high, medium, and low reasoning modes, with the controllable thinking length. Experiment results show that MetaStone-S1 achieves comparable performance to o3-mini’s series with only 32B parameters.
@@ 1.1 Contributions
Reflective generative form: A unified interface for policy model and process reward model
  • We share the backbone parameters between the policy model and SPRM, and design two task-specific heads for token prediction and process scoring respectively (shown in Fig.2). As the policy model has been trained on T-level data during the pre-training stage, the unified interface provides a stable cold-start weight for SPRM, and provides a high way to share knowledge between the two models.
  • The unified interface enables the single network with both reasoning and scoring capabilities. Benefiting from the concise design of the SPRM head, we significantly reduce the extra parameters of the reward model by at least 99%.
  • The unified interface is naturally suitable for TTS. We show a scaling law that reveals the relationship between total thinking computation and TTS performance. Finally, MetaStone-S1 achieves comparable performance as the o3-mini’s series with only 32B parameter size.
Self-superviesd process reward model: Aligning the label of PRM to the outcome reward
[…]
  • The SPRM has a large potential for scaling-up: 1)We find the aha-moment of the SPRM, where the model significantly improves the discrimination between positive and negative thinking processes at a certain checkpoint. 2) SPRM has strong robustness, which is trained on only mathematical data and provides significant improvement in mathematical, coding, and scientific performance.
@@ 2.1 Test-Time Scaling
Test-Time Scaling (TTS) is a technique that leverages additional computational resources at inference time to tackle challenging problems. With the remarkable performance improvements demonstrated by OpenAI o1 (Jaech et al., 2024), TTS has become a research hot spot for enhancing the reasoning capabilities of LLMs. TTS can be divided into two categories: internal TTS and external TTS. Internal TTS introduces the long Chain-of-Thought (CoT) to generate answers based on the detailed reasoning process. OpenAI o1(Jaech et al., 2024) and DeepSeek R1(Guo et al., 2025) introduce a thinking process to plan the solution and guide the final answer. Jin et al. (2024); Yeo et al. (2025) have shown that long CoT can help models correct mistakes by themselves and decompose complex problems more effectively, thereby improving performance. DeepScaler(Luo et al., 2025) demonstrates that by carefully extending the context length during training, only a 1.5B-parameter model can surpass the O1-Preview. However, Chen et al. (2024b, a) have highlighted the risk of overthinking, where excessively long reasoning trajectories may lead to performance degradation. On the other hand, external TTS scaling up inference through search-based strategies and auxiliary reward models. A common approach is the Best-of-N strategy (Lightman et al., 2023; Brown et al., 2024; Wang et al., 2023), which generates multiple candidates and selects the best one based on scores from the pretrained reward model. Moreover, fine-grained methods have also been explored, including, such as Beam Search (Liu et al., 2025; Snell et al., 2024), Diverse Verifier Tree Search (Beeching et al., ) and Monte Carlo Tree Search (MCTS) (Zhang et al., 2024; Guan et al., 2025; Luo et al., 2024). These methods search at the step level and utilize Process Reward Models (PRMs) to guide the reasoning trajectory step-by-step. Beyond search strategies, recent work emphasizes that the quality of the reward model is a crucial factor in external TTS (Guan et al., 2025). A straightforward and effective way to enhance a model’s reasoning ability is to develop a high-quality reward model.
@@ 2.2 Process Reward Model
Process Reward Models (PRMs) focus on evaluating LLMs at the step level. Lightman et al. (2023) unveil that this fine-grained guidance can lead to better TTS performance compared with the global-level Outcome Reward Model (ORM). However, accurately identifying logical errors in LLM outputs remains challenging, PRMs require high-quality task-specific annotated data for training. To this end, recent works Wang et al. (2023) leverage Monte Carlo estimation to automatically assign step-level scores using only the final answers as supervision. Zhang et al. (2024); Guan et al. (2025) iteratively synthesizes data by MCTS and fine-tuning both LLMs and PRMs, improving performance across both models. Tan et al. (2025) follow the LLM-as-a-judge method and introduce a new LLM to annotate the reward of each step. Nonetheless, Zhang et al. (2025) point out that labels generated by Monte Carlo estimation can be noisy, as incorrect reasoning processes may still yield correct final answers. They further propose a hybrid approach that combines both Monte Carlo estimation with the LLM-as-a-judge.
@@
Despite these advances, existing PRMs still suffer from several challenges. First, the PRMs are trained on a new large-scale LLM model, resulting in significant training and inference costs. Second, most PRM training methods typically follow an off-policy strategy, which limits their ability to directly discriminate outputs generated by the target LLM. The unseen distribution during inference time may further degrade performance. To address these issues, we propose a Reflective Generative Form, which shares most parameters with the target LLM and supports on-policy optimization, enabling more efficient and aligned training.
@@
3 Approach
3.1 Overview
Recent Test-Time Scaling (TTS) approaches heavily rely on additional large-scale pretrained reward models (e.g., 7B-72B LLM-based PRMs Liu et al. (2025)), leading to a complex training pipeline and huge extra inference costs. In this work, we propose a Reflective Generative Form, which unifies the reasoning and evaluation processes in a single model. Specifically, a Self-supervised Process Reward Model (SPRM) is introduced, which shares most of its parameters and forward computation with the policy model. During training, we design a Self-supervised Process Reward loss to optimize SPRM only using final answer. The unified model structure and label format allow both reasoning and evaluation tasks to be trained simultaneously in a fully end-to-end manner, eliminating the need for expensive process reward labeling and complex PRM-specific training. At inference time, the policy model and SPRM share the feature extraction parameters, enabling parallel generation of solutions and corresponding evaluation scores. Compared to existing methods with separate PRMs, this design significantly improves TTS efficiency.
In the following sections, we first describe the Reflective Generative Form with SPRM. Then, we introduce the optimization process and present the Self-supervised Process Reward Loss (SPR Loss) for SPRM. Finally, we show the TTS process of Reflective Generative models.
3.2 Reflective Generative Form
The Self-supervised Process Reward Model (SPRM) is proposed to train the policy model to distinguish the solution correctness by itself. The policy model is based on the reasoning LLMs with the thinking process (denoted by the ’<think>’ and ’</think>’ tokens). The joint training framework of the policy model with SPRM is shown in Fig.2(a). In contrast to other PRM methods, SPRM does not need to add additional special tokens for segmenting steps or fine-tune the LLM for step-level output format. It can be simply cold-started from any open-source reasoning model.
Step segmentation.
@@ 3.2 Reflective Generative Form
Process Score Prediction.
A lightweight binary classifier (named SPRM head) is used to predict the score for each step. The SPRM head consists of two linear layers and a dropout layer. Since step-tokens represent the end of each step, we extract features from the second-to-last layer of the model at all step-token positions and feed them into the SPRM head to predict the corresponding process scores. During inference, given a thinking process with multiple steps, we calculate the geometric mean score of all steps as the final score for it, which is formulated as follows:
@@
where $n$ is the number of steps in the given process, and $f_{token_{i}}$ denotes the feature of the $i$-th step-token from the policy model.
With SPRM, the policy model and the reward model are unified into a single interface, allowing them to be trained simultaneously. This combination presents a straightforward and simple training pipeline for on-policy PRM learning, where both models continuously refine their parameters through shared experience, ensuring consistent improvement in the quality of the generated solutions.
3.3 Self-supervised Optimization
For optimizing the policy model, we adopt Group Relative Policy Optimization (GRPO) following Shao et al. (2024). To optimize the SPRM head, we propose a Self-supervised Process Reward Loss (SPR Loss), which enables learning process discrimination ability only from final answer labels. The SPR Loss is formulated as follows,
@@ 3.3 Self-supervised Optimization
where $i$ denotes the step-tokens, $w_{i}$ is a token-level weight, $\text{Score}_{i}$ is the corresponding process score predictions, and $y_{i}$ denotes whether the final prediction is correct. In Eq.2, the process score is optimized based on the correctness of the final answer. However, since a correct final answer may include incorrect intermediate steps and vice versa (Lightman et al., 2023), we introduce the self-supervised dynamic weight $w_{i}$ to mitigate supervision noise. Specifically, we use the SPRM head’s own prediction on each step as the pseudo label and set $w_{i}=1$ only if the pseudo label is consistent with the final answer’s correctness. This dynamic filtering allows the model to avoid noisy samples and focus on the most representative steps of correct and incorrect solutions. Thus, by enlarging the score gap between correct and incorrect steps, SPRM can progressively learn the process evaluation ability with only final annotation.
3.4 Inference with Test-Time Scaling
In the inference stage, the SPRM can guide the thinking process of the policy model through TTS. The total inference stage of SPRM divides into three steps(shown in Fig.2(b)): (1) For the given question, the policy model first samples $k$ candidate thinking process: $think_{1},think_{2},\dots,think_{k}$. (2) The SPRM evaluates the steps in each process and obtains the final score by the geometric mean score of all steps in the corresponding process: $S_{1},S_{2},\dots,S_{k}$. (3) The process with the highest final score is chosen and guides the policy model to answer the question (Eq.3).
$\text{Answer}=\text{LLM}(think_{i^{*}}),\ \text{where}\ i^{*}=argmax(S_{1},S_{2},\dots,S_{k})$ (3)

### v2

@@ Abstract
We introduce our first reflective generative model MetaStone-S1, which obtains OpenAI o3-mini’s performance via the new Reflective Generative Form. The new form focuses on high-quality reasoning trajectory selection and contains two novelties: 1) A unified interface for policy and process reward model: we share the backbone network and use task-specific heads for reasoning trajectory predicting and scoring respectively, introducing only 53M extra parameters for trajectory scoring. 2) Eliminating the reliance on process-level annotation: we provide a self-supervised process reward model, which can directly learn the high-quality reasoning trajectory selection from the outcome reward. Equipped with the reflective generative form, MetaStone-S1 is naturally suitable for test-time scaling, and we provide three reasoning effort modes (low, medium, and high) based on the controllable thinking length. Experiments demonstrate that our MetaStone-S1 achieves comparable performance to OpenAI o3-mini’s series with only 32B parameter size. To support the research community, we have open-sourced MetaStone-S1 at https://github.com/MetaStone-AI/MetaStone-S1.
@@ 1 Introduction
TTS approaches can be categorized into two types: internal TTS and external TTS. Internal TTS (also called sequential TTS in Zeng et al. (2025)) strategies use CoT for longer thinking processes (Guo et al., 2025; OpenAI, 2024), which benefits from Long-CoT Supervised Fine-Tuning and reinforcement learning. Recent internal TTS methods (Guo et al., 2025) mainly suffer from the false positive reasoning process, as the outcome reward will misclassify the correct answer with incorrect reasoning during the training stage. External TTS (also called parallel TTS in Zeng et al. (2025)) is proposed for selecting the correct reasoning process, which is proved to be more effective in performance boosting compared with outcome reward (Lightman et al., 2023). Prominent external TTS algorithms include Best-of-N sampling, Beam Search, and Diverse Verifier Tree Search, using the verifier such as the Process Reward Model(PRM) to select high-quality reasoning trajectories. For example, Best-of-N firstly generates multiple outputs, and then sequentially uses PRM-based to score and select the best solution. Liu et al. (2025) point out that the best external TTS solution varies from policy model with different parameters. When the parameter of the policy model is less than 32B, the search methods (Beam Search and Diverse Verifier Tree Search) achieve better results than the sampling one(Best-of-N). However, when the parameter size is equal to or greater than 32B, the sampling method can achieve better performance. Based on the above analysis, internal and external TTS are two individual methods and can benefit each other, e.g. using the Best-of-N to boost the Long-CoT model with PRM.
This paper focuses on external TTS and proposes a new Reflective Generative Form for high-quality reasoning trajectory selection. Specially, the proposed new form shares the backbone of the policy model and process reward model, and uses self-supervised training to eliminate the reliance on process-level supervision. Based on the Reflective Generative Form, the proposed MetaStone-S1 contains high, medium, and low reasoning modes with the controllable thinking length. Experiment results show that MetaStone-S1 achieves comparable performance to OpenAI o3-mini’s series with only 32B parameters.
@@ 1.1 Contributions
  • A new Reflective Generative Form: We systematically review the existing Test-Time Scaling (TTS) paradigms, and provide a clear definition of the Reflective Generative Form for high-quality reasoning trajectory selection. The proposed Reflective Generative Form enables a single network to achieve both reasoning trajectory prediction and selection (with Zero process-level annotation).
  • The comprehensive analysis about the aha moment, scaling law and robustness of the new form: We provide both qualitative and quantitative analysis for the aha moment, scaling law, and robustness of the proposed new form. These exhaustive discussions will effectively benefit the community for future research.
  • The State-of-the-art performance: MetaStone-S1 achieves comparable performance as the OpenAI o3-mini’s series with only 32B parameter. Specially, MetaStone-S1-low outperforms OpenAI o3-mini-low in mathematical(AIME24&25), coding(LiveCodeBench) and Chinese reasoning(C-Eval) tasks respectively. MetaStone-S1-medium obtains similar results to OpenAI o3-mini-medium. Finally, MetaStone-S1-high further improves the intelligence ceiling and achieves SOTA results in a series of open-source and closed-source models.
@@ 2.1 Test-Time Scaling
Test-Time Scaling (TTS) is a technique that leverages additional computational resources at inference time to tackle challenging problems. With the remarkable performance improvements demonstrated by OpenAI o1 (Jaech et al., 2024), TTS has become a research hot spot for enhancing the reasoning capabilities of LLMs. TTS can be divided into two categories: internal TTS and external TTS. Internal TTS introduces the long Chain-of-Thought (CoT) to generate answers based on the detailed reasoning process. OpenAI o1(Jaech et al., 2024) and DeepSeek R1(Guo et al., 2025) introduce a thinking process to plan the solution and guide the final answer. Jin et al. (2024); Yeo et al. (2025) have shown that long CoT can help models correct mistakes by themselves and decompose complex problems more effectively, thereby improving performance. DeepScaleR(Luo et al., 2025) demonstrates that by carefully extending the context length during training, only a 1.5B-parameter model can surpass the o1-Preview. However, Chen et al. (2024b, a) have highlighted the risk of overthinking, where excessively long reasoning trajectories may lead to performance degradation. On the other hand, external TTS scales up inference through search-based strategies and auxiliary reward models. A common approach is the Best-of-N strategy (Lightman et al., 2023; Brown et al., 2024; Wang et al., 2023), which generates multiple candidates and selects the best one based on scores from the pretrained reward model. Moreover, fine-grained methods have also been explored, including, such as Beam Search (Liu et al., 2025; Snell et al., 2024), Diverse Verifier Tree Search (Beeching et al., ) and Monte Carlo Tree Search (MCTS) (Zhang et al., 2024; Guan et al., 2025; Luo et al., 2024). These methods search at the step level and utilize Process Reward Models (PRMs) to guide the reasoning trajectory step-by-step. Beyond search strategies, recent work emphasizes that the quality of the reward model is a crucial factor in external TTS (Guan et al., 2025). A straightforward and effective way to enhance a model’s reasoning ability is to develop a high-quality reward model.
@@ 2.2 Process Reward Model
Process Reward Models (PRMs) focus on evaluating LLMs at the step level. Lightman et al. (2023) unveil that this fine-grained guidance can lead to better TTS performance compared with the global-level Outcome Reward Model (ORM). However, accurately identifying logical errors in LLM outputs remains challenging, and PRMs require high-quality task-specific annotated data for training. To this end, recent works Wang et al. (2023) leverage Monte Carlo estimation to automatically assign step-level scores using only the final answers as supervision. Zhang et al. (2024); Guan et al. (2025) iteratively synthesizes data by MCTS and fine-tuning both LLMs and PRMs, improving performance across both models. Tan et al. (2025) follow the LLM-as-a-judge method and introduce a new LLM to annotate the reward of each step. Nonetheless, Zhang et al. (2025) point out that labels generated by Monte Carlo estimation can be noisy, as incorrect reasoning processes may still yield correct final answers. They further propose a hybrid approach that combines both Monte Carlo estimation with the LLM-as-a-judge.
@@
Despite these advances, existing PRMs still suffer from several challenges. First, the PRMs are trained on a new large-scale LLM model, resulting in significant training and inference costs. Second, most PRM training methods typically follow an off-policy strategy, which limits their ability to directly discriminate outputs generated by the target LLM. The unseen distribution during inference time may further degrade performance. To address these issues, we propose a Reflective Generative Form, which shares most parameters between the PRM and the target LLM, and supports on-policy optimization with only outcome rewards, enabling more efficient and aligned training.
3 Problem Formulation
This paper aims to find a high-quality reasoning trajectory more efficiently at inference time based on TTS. We first summarize the general inference forms for standard LLMs (policy models) and existing TTS methods, and then formally define our proposed Reflective Generative Form.
1) Basic LLMs. The model directly generates an answer based on the input query. This basic inference form can be formulated as:
$\text{answer}=LLM_{\text{answer}}(\text{query}).$ (1)
TTS based methods can be categorized into two types: sequential scaling based internal TTS and parallel scaling based external TTS.
2) Internal TTS (e.g. DeepSeek R1(Guo et al., 2025)). The internal TTS first generates a reasoning trajectory by Long-CoT using $LLM_{\text{thinking}}$, and then predicts the final answer based on this trajectory using $LLM_{\text{answer}}$. This procedure can be expressed as:
$\text{answer}=LLM_{\text{answer}}(LLM_{\text{thinking}}(\text{query})).$ (2)
To be specific, recent methods (e.g. DeepSeek R1(Guo et al., 2025)) use the same policy model for both $LLM_{\text{thinking}}$ and $LLM_{\text{answer}}$.
3) External TTS (e.g. Lightman et al. (2023); Liu et al. (2025); Zhang et al. (2024)). Firstly, the Long-CoT generation is extended by generating multiple reasoning trajectories and answers in parallel. Then, a reward model (e.g. PRM) is used to score and select the best result. This inference form can be described as:
$\text{answer}=\underset{i\in[1,k]}{\arg\max}\;LLM_{PRM}\Big{(}[LLM_{\text{answer}}(LLM_{\text{thinking}}(\text{query}))]_{i}\Big{)},$ (3)
where $[*]_{i}$ denotes the $i$-th candidate among $k$ parallel generations.
Though existing external TTS methods have been proven to obtain considerable performance enhancement, they still encounter several problems: (1) Extra Computation: PRM contains individual parameters from the policy model($LLM_{\text{think}}$ and $LLM_{\text{answer}}$), which introduces additional huge computation. (2) Expensive Annotation: It is difficult to obtain the large-scale reasoning trajectory annotations for PRM training.
Reflective Generative Form. To address the extra computation and expensive annotation issues, we propose a new Reflective Generative Form focusing on the efficient and label-free reasoning trajectory selection. The proposed Reflective Generative Form is shown in Eq.4.
$\text{answer}=\underbrace{LLM_{\text{answer}}}_{\text{share backbone}}\Big{(}\underset{i\in[1,k]}{\arg\max}\;\underbrace{LLM_{SPRM}}_{\text{share backbone}}\Big{(}[\underbrace{LLM_{\text{thinking}}}_{\text{share backbone}}(\text{query})]_{i}\Big{)}\Big{)}$ (4)
Firstly, we share the backbone of the policy model and PRM in a single network, which enables reasoning trajectory generation and scoring in a unified interface for parallel prediction. The score measures the quality of each reasoning trajectory, and the trajectory with higher score is selected as the high-quality candidate in TTS. This unified interface is proved to be effective for parameter reduction in our experiments. Secondly, we introduce a novel Self-supervised Process Reward Model (SPRM) to eliminate the reliance on process-level annotation, which can be optimized with only outcome-level annotation in a self-supervised manner. In particular, we only implement the SPRM for the $LLM_{\text{think}}$ selection, which can further improve the inference efficiency during the real implementation.
4 Approach
@@
4.1 Unified Interface in Reflective Generative Form
Our proposed Reflective Generative Form establishes a unified interface for the policy model and the PRM. For the policy model, we employ reasoning LLMs that contain the thinking process in response, delineated by the ’<think>’ and ’</think>’ tokens. For the PRM, we introduce a Self-supervised Process Reward Model (SPRM), which shares the same backbone as the policy model but incorporates an additional lightweight SPRM head. The SPRM head is implemented by a binary classifier consisting of two linear layers and a dropout layer. An overview of the joint framework is illustrated in Fig. 2(a).
Within this unified form, the policy model first generates multiple thinking processes as the reasoning trajectories. Subsequently, the SPRM evaluates each thinking process for reasoning trajectory selection. The evaluation procedure contains two steps: (1) Segmenting the reasoning trajectory into discrete steps and (2) Predicting a trajectory score based on evaluation in each step.
@@ 4.1 Unified Interface in Reflective Generative Form
2. Trajectory Score Prediction. After using step-tokens to mark the end of individual reasoning steps, we evaluate each step based on the representation of the corresponding step-token. Since the representation in the last layer mainly captures the logits prediction for a single token, we use the hidden representations from the second-to-last layer of the policy model to provide richer contextual information of the entire step. These representations are then fed into the SPRM head to predict process scores for each step. The final score for the entire reasoning trajectory is computed as the geometric mean of the individual process scores:
@@
where $n$ denotes the total number of steps, and $f_{\text{token}_{i}}$ is the representation of the $i$-th step-token obtained from the policy model. $\text{Score}_{i}$ is the SPRM’s process score for $i$-th step.
Through this unified interface, a single network can generate reasoning trajectories and score them in parallel, enabling joint training in an end-to-end manner. This design facilitates a straightforward and efficient training pipeline for on-policy PRM learning, where both the policy model and the SPRM continuously refine their parameters from shared experiences, thereby improving the overall quality of the generated trajectories.
4.2 Optimization of Reflective Generative Form
During optimization, we train the policy model and the SPRM head simultaneously. For optimizing the policy model, we adopt Group Relative Policy Optimization (GRPO) following Shao et al. (2024). To optimize the SPRM head, we propose a Self-supervised Process Reward Loss (SPR Loss), which enables learning process discrimination ability only from outcome reward (e.g. final answer correctness). The SPR Loss is formulated as follows,
@@ 4.2 Optimization of Reflective Generative Form
where $i$ denotes the step-tokens, $w_{i}$ is a token-level weight, $\text{Score}_{i}$ is SPRM’s process score on step $i$, and $y_{i}$ denotes whether the final answer from the policy model is correct. In Eq.6, the process score is optimized based on the correctness of the final answer. However, since a correct final answer may include incorrect intermediate steps and vice versa (Lightman et al., 2023), we introduce the self-supervised dynamic weight $w_{i}$ to mitigate supervision noise. Specifically, we use the SPRM head’s own prediction on each step as the pseudo label and set $w_{i}=1$ only if the pseudo label is consistent with the final answer’s correctness. This dynamic filtering allows the model to avoid noisy samples and focus on the most representative steps of correct and incorrect solutions. Thus, by enlarging the score gap between correct and incorrect steps, SPRM can progressively learn the process evaluation ability with only final annotations.
4.3 Inference with Reflective Generative Form
In the inference stage, our Reflective Generative Form is naturally suitable for TTS where the SPRM can provide guidance for selecting the high-quality reasoning trajectory from the policy model. The total inference process divides into three steps(shown in Fig.2(b)): (1) For the given question, the policy model first samples $k$ thinking processes as the candidate reasoning trajectories: $think_{1},think_{2},\dots,think_{k}$. (2) The SPRM evaluates the steps in each process and obtains the final score by the geometric mean of corresponding process scores: $S_{1},S_{2},\dots,S_{k}$. (3) The reasoning trajectory with the highest final score is chosen and guides the policy model to answer the question (Eq.7).
$\text{answer}=LLM_{\text{answer}}(think_{j}),\ \text{where}\ j=argmax(S_{1},S_{2},\dots,S_{k})$ (7)